OpenStudy (anonymous):

2 + 5(3-6)+ 42 - 10 = ? I cant get the answer can someone explain??

OpenStudy (anonymous):

19 order of operations

OpenStudy (anonymous):

2+5(3-6)+42-10 Step 1: Parentheses first 3-6 = -3 2+5(-3) + 42 - 10 Step 2: Multiply 2 + (-15) + 42 - 10 Step 3: Solve 2 - 15 + 42 -10 2- 15 = -13 42 - 10 = 32 -13 + 32 = 19 Answer = 19

OpenStudy (anonymous):

i understand now, thanks for the knowledge